OpenStudy (anonymous):

If (px+q)^2=9x^2+kx+16 what is the value of k?

OpenStudy (anonymous):

We know that, for a quadratic expression to be monomially-factorable, we have: \[ (ax+b)^2=a^2x^2+2abx+b^2 \]So, therefore, since we wish to find \(2ab\) in terms of \(a^2\) and \(b^2\): \[ 2ab=2\sqrt{a^2b^2} \]With this, we solve: \[ k=2ab=2\sqrt{9\cdot16}=2\cdot3\cdot4=24=k \]Et voilá.